A particular company produces widgets. Each widget sells for $9, and the variable cost of producing each units is 40% of the selling price. If the monthly fixed costs incurred by the company are $50,000, what is the break-even point?
Answer by mananth(16946)

let the number of widgets produced be x
..
Sale price = $9
..
Total sales = 9x
..
Cost of production =0.4*9= $3.60
..
50,000+3.60x = 9x
50,000=9x-3.6x
50,000 = 5.4x
/5.4
50,000/5.4
x= 9260 Break even point
...
m.ananth@hotmail.ca